OpenStudy (anonymous):

Multiple and simplify the product (2sqrt{7}-\sqrt{3})^2

OpenStudy (anonymous):

(2 sqrt/(7)-sqrt(3))^2 = 4*7 -(2*2sqrt(7)*sqrt(3))+3 = 31-(4*sqrt(7)*sqrt(3)) ~= 12.67